OpenStudy (anonymous):

if X and Y are two whole numbers and their relation is X-Y =1 ,then if X^2-Y^2=995283457 so what are the values of X and Y?

OpenStudy (anonymous):

does it help to know that \(x^2-y^2=(x+y)(x-y)\)

OpenStudy (anonymous):

in other words, the question is asking this: find two consecutive integers whose sum is \(995,283,457\)

OpenStudy (amistre64):

let x=1+n, let y=n then solve for n in the hyperbola ... so yeah, 2 consecutives